The Earth is approximately a sphere of radius 6.37 x 106 m. a) Calculate the distance from the pole to the equator, measured along the surface of the Earth. b) Calculate the distance from the pole to the equator, measured along a straight line passing through the Earth.
1
Expert's answer
2020-04-20T10:43:54-0400

a)


"d=\frac{\pi}{2}R=\frac{\pi}{2}6.37 \cdot10^6=1.0\cdot10^7m"

b)


"d=\sqrt{2}R=\sqrt{2}\cdot6.37 \cdot10^6=9.0\cdot10^6m"